OpenStudy (anonymous):

Evaluate : f(-3) if f(x)=2|4-x|+6

OpenStudy (anonymous):

put x=-3 what is |4-3| f(-3)=2*1+6 simplify

OpenStudy (anonymous):

f(-3)=8 Divide by three is this right? @giligails

OpenStudy (anonymous):

no it is not division i think it is absolute value absolute value means |-x|=x

OpenStudy (anonymous):

so i think |4-(-3)|=7 2*7=14 14+6=20

OpenStudy (anonymous):

Okay that makes sense